# Molecular sorting on a fluctuating membrane

### v1

@@ Abstract
Introduction. Molecular sorting is a vital process in eukaryotic cells, where proteins and other biomolecules are sorted and encapsulated into lipid vesicles for targeted transport to specific subcellular locations. This distillation process occurs on lipid membranes, such as the plasma membrane [1], endosomes, the Golgi apparatus [2], and the endoplasmic reticulum [3], where biomolecules can bind and diffuse laterally. Due to a variety of direct and indirect interactions, these molecules aggregate into domains with distinct chemical compositions. These domains can induce membrane bending and fission, ultimately forming separated submicron lipid vesicles that are transported to their designated subcellular sites by molecular motors. In this way, lipid membranes act as natural molecular distillers, promoting intracellular order and compartmentalization and counteracting the homogenizing effects of diffusion. Disruption of molecular sorting in living cells is implicated in severe pathologies, including cancer [4, 5]. On the other end of the spectrum, analogous molecular sorting processes are exploited by enveloped viruses, such as HIV, SARS-CoV, and influenza, for their assembly and budding from host cells [6, 7, 8, 9], further underscoring the practical relevance of understanding the physical mechanisms of molecular sorting.
@@
Phase separation in living cells is driven by short-range, weakly polar electrostatic interactions between disordered regions of proteins [15], as well as by reinforcing enzymatic feedback loops, as in diffusion-limited phase separation [16]. On the other hand, it has long been established that protein inclusions in lipid membranes are subject to Casimir-like interactions [17, *GBP93b], a significant class of entropic interactions. However, their role in molecular sorting remains unexplored. These interactions arise from the increased rigidity of the membrane in the presence of embedded protein inclusions, which restricts membrane fluctuations and generates entropic attractive forces. It is known that proteins and lipids involved in the formation of sorting domains increase local membrane rigidity by a factor of 10 to 30 compared to the surrounding membrane [19, 20, 21]. Here, we investigate the role of these Casimir-like interactions in molecular sorting and find that they can significantly enhance the molecular sorting process, especially within the biologically relevant regime of weak short-range interactions.
Phenomenological Theory. Building on our previous work, we investigate the role of the lipid membrane as a distiller of molecular species [10, 11, 12]. In this scenario, molecules are randomly inserted into the membrane, diffuse laterally, and aggregate into sorting domains due to short-range attractive forces. The sorting domains grow by adsorbing molecules from the surrounding “gas” of freely diffusing molecules. Domains of size $R$ larger than a critical value $R_{\mathrm{c}}$ grow irreversibly through the absorption of single molecules diffusing towards them [22, 23, 11]. The growth rate is determined by the net flux $\Phi$ of molecules towards a domain, which in turn is proportional to the molecular density difference $\Delta n=n_{L}-n_{R}$ between distant regions and regions adjacent to the domain boundaries [10]. Domains that reach a characteristic size $R_{\mathrm{E}}$ are ultimately removed from the membrane through the formation of small, separate lipid vesicles [10].
@@
We have shown in Ref. 10 that an optimal sorting regime is achieved for an intermediate strength of short-range attractive forces. When the tendency to aggregate is too strong, a proliferation of slowly growing sorting domains occurs, leading to molecular crowding and decreased sorting efficiency [10, 12]. In the optimal sorting regime, there exists a specific density $N_{\mathrm{d}}$ of sorting domains, resulting in minimal average molecular density [10]. For absorbing domains, the average residence time $T$ of a molecule of linear size $a$ in the membrane system is the sum of the average time $T_{\mathrm{f}}$ required for the molecule to reach a sorting domain by free diffusion and be absorbed, and the average time $T_{\mathrm{d}}$ spent inside the domain until the extraction event. The two contributions can be estimated as [10]
@@
where $A$ is a dimensionless, increasing function of the relative rigidity $\alpha=\kappa_{1}/\kappa_{0}$ (see Supplemental Material). Eq. (4) implies that $U\sim A\,k_{\mathrm{B}}T$ near the surface of a domain. On the other hand, the interaction potential between two inclusions mediated by the membrane fluctuations decays as $r^{-4}$ for distances $r$ much larger than their sizes [17].
@@
where $R$ is the radius of the domain and $n(R)$ is the molecular density near the domain boundary. The potential $U$, induced by membrane fluctuations, is of the order of $k_{\mathrm{B}}T$ when $r\sim R$ and tends to zero as $r$ grows (see Supplemental Material). Thus, for $r\gg R$, the solution (6) reduces to
@@
Numerical results. To validate our theoretical predictions, we implemented a numerical scheme that generalizes the lattice-gas model of molecular sorting introduced in Ref. 10. This scheme shares several features with the approach used in Ref. 41 to investigate the phase separation of rigid inclusions in fluid membranes close to thermodynamic equilibrium, although we are studying here an out-of-equilibrium state. We consider a fluctuating membrane described by a discretized version of Helfrich Hamiltonian, on which inserted molecules laterally diffuse and aggregate. The system is driven out of equilibrium by an incoming flux of molecules, which are randomly attached at empty membrane sites with a rate $\phi$ per unit area, and is maintained in a statistical stationary state by the instantaneous removal of connected molecular domains that reach the threshold number of molecules $N_{\mathrm{E}}$. Consistently with our theoretical approach, simulations are performed in the adiabatic regime.
@@
Membrane configurations are sampled using a Monte Carlo algorithm. After each Monte Carlo sweep (MCS), steps involving molecule insertion, diffusion, and the extraction of domains of size $\geq N_{\mathrm{E}}$ are performed. One MCS is taken as the time unit. The rate of molecule insertion per empty site is denoted by $k_{\mathrm{I}}$. The diffusion rate $k_{\mathrm{D}}$ of free molecules is measured as the ratio of accepted diffusive jumps during one MCS (see Appendix for additional details). Simulations are performed with the realistic parameter values $\kappa_{0}=10\,k_{\mathrm{B}}T$, $N_{\mathrm{E}}=25$ [42, 43, 10, 11, 12], while $k_{I}$ and $k_{D}$ are kept much smaller than $1$ in inverse MCS units, to ensure proper sampling of membrane configurations within the adiabatic regime.
@@
We investigated the behavior of the density $\bar{\rho}$ as a function of the short-range interaction $W$ and molecular rigidity $\kappa_{1}$. In Fig. 1, the resulting stationary densities are plotted as functions of the short-range interaction strength $W$ for the fixed dimensionless flux $\phi/k_{\mathrm{D}}=10^{-5}$ (see Appendix), with varying $\alpha=\kappa_{1}/\kappa_{0}$.
@@
Conclusions. The lipid membranes of endosomes, the Golgi apparatus, the endoplasmic reticulum, and the plasma membrane play a fundamental role in sorting and distilling vital molecular factors, acting as a natural realization of Szilard’s model of classical nucleation theory [23]. These delicate structures are inherently subject to thermally induced fluctuations. Previous studies have shown that such fluctuations significantly contribute to the phase separation of rigid membrane inclusions close to thermodynamic equilibrium [41]. Our analysis extends these findings to the out-of-equilibrium scenario of molecular sorting, demonstrating that membrane-mediated interactions can strongly enhance the molecular distillation of rigid inclusions, particularly, in the biologically relevant regimes where short-range intermolecular attractive forces are relatively weak. Our analysis suggests that thanks to membrane-mediated interactions, rigid biomolecules can be sorted with high efficiency, despite their low-affinity interactions. Notably, this effect, potentially crucial for biological systems, is observed in our numerical simulations well below the threshold where phase separation occurs close to equilibrium [41]. This suggests an important distinction between classical quasi-equilibrium phase separation processes and the role phase separation plays in out-of-equilibrium biological systems.
Molecular inclusions interact with the surrounding membrane due to both their rigidity and, possibly, non-zero intrinsic curvature [17]. In this study, we have focused on the impact of rigidity on the molecular sorting process. In future work, we plan to investigate the complex interplay between rigidity and intrinsic curvature.
Our findings suggest that a key parameter controlling molecular sorting efficiency is the relative rigidity of the membrane and supermolecular domains, and that higher efficiency is achieved at intermediate values of this relative rigidity. These theoretical predictions may inspire future experimental investigations in real biological cells or in artificially prepared membranes.
@@ Acknowledgements.
Acknowledgments. AG would like to thank Guido Serini for many fruitful discussions. Numerical calculations have been made possible through a CINECA-INFN agreement providing access to computational resources at CINECA.
Appendix A Appendix – Simulation protocol

### v2

@@ Abstract
I Introduction
Molecular sorting is a vital process in eukaryotic cells, where proteins and other biomolecules are sorted and encapsulated into lipid vesicles for targeted transport to specific subcellular locations. This distillation process occurs on lipid membranes, such as the plasma membrane [1], endosomes, the Golgi apparatus [2], and the endoplasmic reticulum [3], where biomolecules can bind and diffuse laterally. Due to a variety of direct and indirect interactions, these molecules aggregate into domains with distinct chemical compositions. These domains can induce membrane bending and fission [4, 5, 6, 7], ultimately forming separated submicron lipid vesicles that are transported to their designated subcellular sites by molecular motors. In this way, lipid membranes act as natural molecular distillers, promoting intracellular order and compartmentalization and counteracting the homogenizing effects of diffusion. Disruption of molecular sorting in living cells is implicated in severe pathologies, including cancer [8, 9]. On the other end of the spectrum, analogous molecular sorting processes are exploited by enveloped viruses, such as HIV, SARS-CoV, and influenza, for their assembly and budding from host cells [10, 11, 12, 13], further underscoring the practical relevance of understanding the physical mechanisms of molecular sorting.
@@ I Introduction
Phase separation is emerging as one of the main ordering processes in living cells [19, 20, 21], and various mechanisms have been proposed as its drivers. Among them, weakly polar electrostatic interactions between disordered regions of proteins [22], active processes, as in diffusion-limited phase separation, mass-conserved reaction-diffusion systems and active emulsions [23, 24, 25, 26, 27, 28, 29], and segregating kinetic effects [30]. On the other hand, it has long been established that protein inclusions in lipid membranes are subject to Casimir-like interactions [31, *GBP93b, 33], a significant class of entropic interactions. However, their role in molecular sorting remains unexplored. These interactions arise from the increased rigidity of the membrane in the presence of embedded protein inclusions, which restricts membrane fluctuations and generates entropic attractive forces. It is known that proteins and lipids involved in the formation of sorting domains increase local membrane rigidity by a factor of 10 to 30 compared to the surrounding membrane [34, 35, 36]. Here, we investigate the role of these Casimir-like interactions in molecular sorting and find that they can significantly enhance the molecular sorting process, especially within the biologically relevant regime of weak short-range interactions.
II Phenomenological Theory
Building on our previous work, we investigate the role of the lipid membrane as a distiller of molecular species [14, 15, 16]. In this scenario, molecules are randomly inserted into the membrane, diffuse laterally, and aggregate into sorting domains due to short-range attractive forces. The sorting domains grow by adsorbing molecules from the surrounding “gas” of freely diffusing molecules. Domains of size $R$ larger than a critical value $R_{\mathrm{c}}$ grow irreversibly through the absorption of single molecules diffusing towards them [37, 38, 15]. The growth rate is determined by the net flux $\Phi$ of molecules towards a domain, which in turn is proportional to the molecular density difference $\Delta n=n_{L}-n_{R}$ between distant regions and regions adjacent to the domain boundaries [14]. Domains that reach a characteristic size $R_{\mathrm{E}}$ are ultimately removed from the membrane through the formation of small, separate lipid vesicles [14].
@@ II Phenomenological Theory
We have shown in Ref. 14 that an optimal sorting regime is achieved for an intermediate strength of short-range attractive forces. When the tendency to aggregate is too strong, a proliferation of slowly growing sorting domains occurs, leading to molecular crowding and decreased sorting efficiency [14, 16]. In the optimal sorting regime, there exists a specific density $N_{\mathrm{d}}$ of sorting domains, resulting in minimal average molecular density [14]. For absorbing domains, the average residence time $T$ of a molecule of linear size $a$ in the membrane system is the sum of the average time $T_{\mathrm{f}}$ required for the molecule to reach a sorting domain by free diffusion and be absorbed, and the average time $T_{\mathrm{d}}$ spent inside the domain until the extraction event. The two contributions can be estimated as [14]
@@
where $A$ is a dimensionless, increasing function of the relative rigidity $\alpha=\kappa_{1}/\kappa_{0}$ (see Appendix A). Eq. (23) implies that $U\sim A\,k_{\mathrm{B}}T$ near the surface of a domain. On the other hand, the interaction potential between two inclusions mediated by the membrane fluctuations decays as $r^{-4}$ for distances $r$ much larger than their sizes [31].
@@
where $R$ is the radius of the domain and $n(R)$ is the molecular density near the domain boundary. The potential $U$, induced by membrane fluctuations, is of the order of $k_{\mathrm{B}}T$ when $r\sim R$ and tends to zero as $r$ grows (see Appendix A). Thus, for $r\gg R$, the solution (6) reduces to
@@
III Numerical results
To validate our theoretical predictions, we implemented a numerical scheme that generalizes the lattice-gas model of molecular sorting introduced in Ref. 14. This scheme shares several features with the approach used in Ref. 56 to investigate the phase separation of rigid inclusions in fluid membranes close to thermodynamic equilibrium, although we are studying here an out-of-equilibrium state. We consider a fluctuating membrane described by a discretized version of Helfrich Hamiltonian, on which inserted molecules laterally diffuse and aggregate. The system is driven out of equilibrium by an incoming flux of molecules, which are randomly attached at empty membrane sites with a rate $\phi$ per unit area, and is maintained in a statistical stationary state by the instantaneous removal of connected molecular domains that reach the threshold number of molecules $N_{\mathrm{E}}$. Consistently with our theoretical approach, simulations are performed in the adiabatic regime.
@@ III Numerical results
Membrane configurations are sampled using a Monte Carlo algorithm. After each Monte Carlo sweep (MCS), steps involving molecule insertion, diffusion, and the extraction of domains of size $\geq N_{\mathrm{E}}$ are performed. One MCS is taken as the time unit.
The rate of molecule insertion per empty site is denoted by $k_{\mathrm{I}}$. The diffusion rate $k_{\mathrm{D}}$ of free molecules is measured as the ratio of accepted diffusive jumps during one MCS (see Appendix B for additional details).
Figure 3: Characterization of the sorting process in the statistically steady state in terms of three key observables, measured from numerical simulations as functions of the short-range interaction strength $W$, for varying relative rigidities $\alpha=\kappa_{1}/\kappa_{0}$: (a) the critical radius $R_{\mathrm{c}}$ (estimated using the method described in Ref. 15); (b) the number density $N_{\mathrm{d}}$ of supercritical domains; and (c) the average density of isolated molecules $\bar{n}$. Due to the logarithmic profile of the molecular density around sorting domains, the average density $\bar{n}$ is close to $n_{L}$.
Simulations are performed with the realistic parameter values $\kappa_{0}=10\,k_{\mathrm{B}}T$, $N_{\mathrm{E}}=25$ [57, 58, 14, 15, 16], while $k_{I}$ and $k_{D}$ are kept much smaller than $1$ in inverse MCS units, to ensure proper sampling of membrane configurations within the adiabatic regime.
@@
We investigated the behavior of the density $\bar{\rho}$ as a function of the short-range interaction $W$ and molecular rigidity $\kappa_{1}$. In Fig. 2, the resulting stationary densities are plotted as functions of the short-range interaction strength $W$ for the fixed dimensionless flux $\phi/k_{\mathrm{D}}=10^{-5}$ (see Appendix B), with varying $\alpha=\kappa_{1}/\kappa_{0}$.
@@
IV Conclusions
The lipid membranes of endosomes, the Golgi apparatus, the endoplasmic reticulum, and the plasma membrane play a fundamental role in sorting and distilling vital molecular factors, acting as a natural realization of Szilard’s model of classical nucleation theory [38]. These delicate structures are inherently subject to thermally induced fluctuations. Previous studies have shown that such fluctuations significantly contribute to the phase separation of rigid membrane inclusions close to thermodynamic equilibrium [56]. Our analysis extends these findings to the out-of-equilibrium scenario of molecular sorting, demonstrating that membrane-mediated interactions can strongly enhance the molecular distillation of rigid inclusions, particularly, in the biologically relevant regimes where short-range intermolecular attractive forces are relatively weak. Our analysis suggests that thanks to membrane-mediated interactions, rigid biomolecules can be sorted with high efficiency, despite their low-affinity interactions. Notably, this effect, potentially crucial for biological systems, is observed in our numerical simulations well below the threshold where phase separation occurs close to equilibrium [56]. This suggests an important distinction between classical quasi-equilibrium phase separation processes and the role phase separation plays in out-of-equilibrium biological systems.
Molecular inclusions interact with the surrounding membrane due to both their rigidity and, possibly, non-zero intrinsic curvature [31, 60] In this study, we have focused on the impact of rigidity on the molecular sorting process. In future work, we plan to investigate the complex interplay between rigidity and intrinsic curvature.
Our findings suggest that a key parameter controlling molecular sorting efficiency is the relative rigidity of the membrane and supermolecular domains, and that higher efficiency is achieved at intermediate values of this relative rigidity. These predictions may inspire future experimental investigations in real biological cells or in artificially prepared membranes.
@@ Acknowledgements.
AG would like to thank Guido Serini for many fruitful discussions. Numerical calculations have been made possible through a CINECA-INFN agreement providing access to computational resources at CINECA.
Appendix A Interaction of a molecule with a domain
In this section, we analyze the Casimir interaction between a circular domain of radius $R$ and a single molecule of radius $a\ll R$, positioned at a distance $x\gg a$ from it. We will calculate the interaction potential between the molecule and the domain.
In the absence of overhangs, the membrane can be parameterized in the Monge gauge [61], where each point on the membrane is defined by its displacement $u(\bm{r})=u(x,y)$ in the direction perpendicular to a reference plane $\mathcal{S}$. To second order in $u$, the Helfrich Hamiltonian, which provides the elastic energy of the deformed membrane, reads
$\mathcal{H}$ $=\int_{\mathcal{S}}\mathrm{d}x\,\mathrm{d}y\,\left\{\frac{\kappa}{2}(\nabla^{2}u)^{2}+\bar{\kappa}[\partial_{x}^{2}u\,\partial_{y}^{2}u-(\partial_{x}\partial_{y}u)^{2}]\right\},$ (12)
Here $\kappa$ and $\bar{\kappa}$ are bending and Gaussian rigidities, determined by an internal structure of the membrane. A surface-tension contribution to the energy could also be included, but it is assumed to be negligible and will not be taken into account.
Here we consider the interaction of a single molecule with a circular domain of molecules inserted into the membrane. When the molecule is positioned at the point $\bm{r}=(x,y)$, the interaction potential of the molecule with the domain is
$U$ $=$ $\phantom{+}B(\partial_{x}^{2}\partial_{x^{\prime}}^{2}\mathcal{G}|_{x=x^{\prime},y=y^{\prime}}+2\partial_{x}^{2}\partial_{y^{\prime}}^{2}\mathcal{G}|_{x=x^{\prime},y=y^{\prime}}$ (13)
$+\partial_{y}^{2}\partial_{y^{\prime}}^{2}\mathcal{G}|_{x=x^{\prime},y=y^{\prime}})+D(\partial_{x}^{2}\partial_{y^{\prime}}^{2}\mathcal{G}|_{x=x^{\prime},y=y^{\prime}}$
$-\partial_{x}\partial_{y}\partial_{x^{\prime}}\partial_{y^{\prime}}\mathcal{G}|_{x=x^{\prime},y=y^{\prime}})$
where ${\mathcal{G}}(\bm{r},\bm{r}^{\prime})$ is the contribution to the pair correlation function $\langle u(\bm{r})u(\bm{r}^{\prime})\rangle$ from the membrane displacement induced by the domain. The factors $B,D$ in Eq. (13) are introduced via the phenomenological coupling energy of the molecule with the membrane, when the former is treated as a point-like object:
$\delta\mathcal{H}=B(\nabla^{2}u)^{2}+D[\partial_{x}^{2}u\partial_{y}^{2}u-(\partial_{x}\partial_{y}u)^{2}]$ (14)
where the derivatives are evaluated at the position of the molecule. This expression is valid for fluctuations of $u$ on scales much larger than $a$. The factors $B$ and $D$ are functions of the rigidity and size of the molecule. We will make use of the fact that their expression for a disc of radius $a$ and rigidity $\kappa=\kappa_{1},\bar{\kappa}=-\kappa_{1}$, inserted in a membrane of rigidity $\kappa=\kappa_{0},\bar{\kappa}=-\kappa_{0}$ is 111 The detailed derivation of these expressions will be published in a future work. :
$B$ $=$ $\pi a^{2}\kappa_{0}(\kappa_{1}-\kappa_{0})\left(\frac{1}{(\kappa_{1}+\kappa_{0})}+\frac{1}{\kappa_{1}+3\kappa_{0}}\right)$
$D$ $=$ $-\pi a^{2}\frac{4(\kappa_{1}-\kappa_{0})\kappa_{0}}{\kappa_{1}+3\kappa_{0}}.$ (15)
If the separation between the molecule and the domain boundary is much smaller than the domain size $R$, the boundary can be approximated as a straight line. Therefore, we assume that the domain occupies the half-plane $x<0$. We also consider that the domain and the bulk membrane have different bending and Gaussian rigidities, $\kappa_{1},\bar{\kappa}_{1}$ and $\kappa_{0},\bar{\kappa}_{0}$. respectively. The Hamiltonian of the system is then given by
[…]
where $\mathcal{D}_{1}$ is the left half-plane ($x<0$) and $\mathcal{D}_{2}$ is the right half-plane ($x>0$).
Using linear response theory, we can derive an equation for the pair correlation function $G=\langle u(\bm{r})u(\bm{r}^{\prime})\rangle$, entering Eq. (13). It is important to note here that, due to the system’s homogeneity in the $y$ direction and its invariance under reflection $y\to-y$, $G$ is a function of $|y-y^{\prime}|$. The resulting equations read
$\nabla^{4}G$ $=$ $\frac{k_{\mathrm{B}}T}{\kappa_{1}}\delta(x-x^{\prime})\delta(y-y^{\prime})\qquad x<0$
$\nabla^{4}G$ $=$ $\frac{k_{\mathrm{B}}T}{\kappa_{0}}\delta(x-x^{\prime})\delta(y-y^{\prime})\qquad x>0$ (17)
with boundary conditions
$\partial_{x}(\kappa_{1}\nabla^{2}-\bar{\kappa}_{1}\partial_{y}^{2})G|_{x=0^{-}}$ $=$ $\partial_{x}(\kappa_{0}\nabla^{2}-\bar{\kappa}_{0}\partial_{y}^{2})G|_{x=0^{+}}$
$(\kappa_{1}\nabla^{2}+\bar{\kappa}_{1}\partial_{y}^{2})G|_{x=0^{-}}$ $=$ $(\kappa_{0}\nabla^{2}+\bar{\kappa}_{0}\partial_{y}^{2})G|_{x=0^{+}}$ (18)
Observe that, due to the inhomogeneity of the Gaussian rigidity, the topological term involving Gaussian curvature in the Hamiltonian cannot be neglected. This term contributes to the boundary conditions (18) for the correlation function.
Due to translation invariance along the $y$ direction, it is convenient to make use of the Fourier transform
$\hat{G}(x,x^{\prime},q)=\int_{-\infty}^{+\infty}\mathrm{d}y\,\exp[iq(y-y^{\prime})]G(x,x^{\prime},y-y^{\prime}),$ (19)
which is an even function of $q$. The solutions to Eqs. (17) and (18) for $q>0$ are
$\hat{G}(x,x^{\prime},q)=(A_{0}+A_{1}x)e^{qx}+\frac{k_{\mathrm{B}}T}{4q^{3}\kappa_{1}}(1+q|x-x^{\prime}|)e^{-q|x-x^{\prime}|}$ (20)
for $x<0$, and
$\hat{G}(x,x^{\prime},q)=(B_{0}+B_{1}x)e^{-qx}+\frac{k_{\mathrm{B}}T}{4q^{3}\kappa_{0}}(1+q|x-x^{\prime}|)e^{-q|x-x^{\prime}|}$ (21)
for $x>0$. The factors $A_{0},A_{1},B_{0},B_{1}$ must be determined from the continuity of $\hat{G}$ and its derivative $\partial_{x}\hat{G}$ at $x=0$ and from the boundary conditions (18), where $\partial_{r}^{2}\to-q^{2}$, $\nabla^{2}\to\partial_{x}^{2}-q^{2}$. Assuming $\bar{\kappa}_{0}=-\kappa_{0}$ and $\bar{\kappa}_{1}=-\kappa_{1}$, the correlation function for $x,x^{\prime}>0$ is
$\hat{G}(x,x^{\prime},q)$ $=$ $\frac{k_{\mathrm{B}}T}{4q^{3}\kappa_{0}}\left[(1+q|x-x^{\prime}|)e^{-q|x-x^{\prime}|}-\frac{e^{-q(x+x^{\prime})}(\kappa_{1}-\kappa_{0})((3\kappa_{1}+\kappa_{0})q(x+x^{\prime}+2qxx^{\prime})+3\kappa_{1}+5\kappa_{0})}{(3\kappa_{1}+\kappa_{0})(\kappa_{1}+3\kappa_{0})}\right].$ (22)
The second term in the square brackets determines the contribution $\mathcal{G}$ to the correlation function induced by the domain.
In accordance with Eqs. (13,15,22) the interaction energy of the molecule with the domain is
$U(x)=-A\,k_{\mathrm{B}}T\,\frac{a^{2}}{x^{2}}$ (23)
where, letting $\alpha=\kappa_{1}/\kappa_{0}$,
$A=\frac{(\alpha-1)^{2}(3\alpha+5)(5\alpha+3)}{4(\alpha+1)(\alpha+3)^{2}(3\alpha+1)}$ (24)
The factor $A$ is a monotonically increasing function of $\alpha$ for $\alpha>1$.
At large separations between the molecule and the domain, the size $R$ of the domain becomes a relevant scale, and its boundary can no longer be treated as an infinite wall. In this case, the interaction can be evaluated as [63]
$U(x)=k_{\mathrm{B}}T\frac{BD_{R}+B_{R}D}{2\pi^{2}\kappa_{0}^{2}x^{4}}=-\tilde{A}k_{\mathrm{B}}T\frac{a^{2}R^{2}}{x^{4}},$ (25)
where
$\tilde{A}=\frac{2(\alpha-1)^{2}(3\alpha+5)}{(\alpha+1)(\alpha+3)^{2}}.$ (26)
Note that, by taking the appropriate limits, this expression reproduces previous analytical results found in the literature [31, 52].
When considering a single molecule diffusing in the vicinity of a sorting domain, one of the two regimes in Eq. (23) and Eq. (25) should be considered depending on the distance. A convenient interpolation formula for the membrane-mediated interaction energy between a molecule and a sorting domain of radius $R$, valid across different asymptotic regimes, is given by the simplest two-point Padé approximant [64]
$U(r)$ $=-\hskip 1.69998ptk_{\mathrm{B}}T\,\frac{R^{2}}{r^{2}}\left[\frac{A\,a^{2}}{(r-R)^{2}+a^{2}}+(\tilde{A}-A)\frac{a^{2}}{r^{2}}\right]$ (27)
where $r=x+R$ is the distance from the molecule to the center of the domain. This reduces to Eq. 25 when $r\gg R$, $r\gg a$, and to Eq. 23 in the limit $r\sim R$ and $r-R\gg a$, while also avoiding the unphysical singularity at $x=0$.
Appendix B Simulation protocol